OpenStudy (anonymous):

x^2-20x=10 complete the square

OpenStudy (anonymous):

x^2-20x-10=0 use the quadratic formula http://www.mathsisfun.com/quadratic-equation-solver.html

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

It doesnt let me use the minus sign.

OpenStudy (anonymous):

Complete the square means to complete the square. Like this: x^2 - 20x is the square of what. Well, if we add + 100 x^2 - 20x + 100 = (x - 10)(x - 10). So, we have x^2 - 20x - 10 = x^2 - 20x - 10 + 100 - 100 = (x-10)^2 - 110 x = sqrt(110) + 10 The quadratic gives you: 20 +/- sqrt (400 + 40) / 2 = 10 + sqrt(110) So it checks.

OpenStudy (anonymous):

Thanks.